'If I use 0.0147 kg of glucose (C6H12O6) and 0.150 L of water to make a solution (density of water = 1.00 g/mL; Kb=0.512oC/m.), how much have I changed the boiling point of the water?



A.0.279°C
B.0.502°C
C.0.0418°C
D.5.22°C
E.none of the above

'You need to make 6.50 L of a sulfuric acid solution that has a concentration of 1.25 M. If you are going to use 9.0 M as your stock solution to make the dilution, how much of the stock solution should you use?



A.1.10 L
B.748 mL
C.1.73 L
D.903 mL
1
Expert's answer
2017-02-06T07:43:34-0500
1. m(H₂O)=₰*V=1g/ml*150ml=150g
n(C₆H₁₂O₆)=m(C₆H₁₂O₆)/M(C₆H₁₂O₆)=14,7g/180g/Mo=0,0817Mol
m´= n(C₆H₁₂O₆)/m(H₂O)=0,0817/0,150=0,54Mol/kg
Δt=K*m´; Δt=0,512*,05447=0,279ᵒC
Answer A

2. During the rule of the cross
9Mol(H₂SO₄) 1,25parts
1,25Mol 7,2parts
1Mol(H₂O) 7,75parts

6,5:7,2=0,903ml
Answer D